OpenStudy (anonymous):

in preparation for the final exam, which counts as two exams, a student has scores of 88,97, and 91. what does the student need on the final exam to have a 90 average?

OpenStudy (anonymous):

(88.97+91+x)/3=90 180-88.97-91=x

OpenStudy (anonymous):

missed out a term mr stats major :P

OpenStudy (anonymous):

oopsies, please correct me, mr. troll :D

OpenStudy (anonymous):

oh wow, I was reading that as 88.97; der... should be (88+97+91+x)/4=90 360-88-97-91=x=84

OpenStudy (anonymous):

no, well I guess you didnt miss out a term, but you read the first term as 88.97 , ill bet ohh I dont know $1000 that they meant it to be two seperate terms now mr stats major, what would be my chances of being correct and taking your money :D

OpenStudy (anonymous):

@elecengineer: lol, you sure it was 8,897?

OpenStudy (anonymous):

ohh dont worry, you are only a dumb stats major with no real use to society , just play around with lots of numbers and probabilities, so boring

OpenStudy (anonymous):

hahahhahah we both got it wrong; the final counted as two exams brosky: (88+97+91+2x)/5=90

OpenStudy (anonymous):

u mean you got it wrong

OpenStudy (anonymous):

and you didn't help ;)